OpenStudy (anonymous):

which function has the following characteristics? vertical asymptote at x=3. and horizontal asymptote at y=2. Domain {x =with a cross through it plus minus 3}

OpenStudy (smored):

\[x \neq \pm3\] that?

OpenStudy (anonymous):

yeah

OpenStudy (smored):

tkhunny shall answer

OpenStudy (anonymous):

do I need to put anything else?

OpenStudy (tkhunny):

You should put the answer choices, perhaps? :-) Start with a bare function and build it a step at a time. \(y = \dfrac{}{}\) vertical asymptote at x=3 \(y = \dfrac{}{x-3}\) horizontal asymptote at y=2 \(y = \dfrac{2x}{x-3}\) Domain {x =with a cross through it plus minus 3} This hint is a little tricky. There is a vertical asymptote at x = 3, so x = 3 is not in the Domain. What's the deal with x = -3? If there is a vertical asymptote at x = -3, there are two necessary adjustments. \(y = \dfrac{2x^{2}}{(x-3)(x+3)}\) If there is a NOT vertical asymptote at x = -3, it's just missing, there are two necessary adjustments. \(y = \dfrac{2x(x+3)}{(x-3)(x+3)}\) This is NOT a unique function. There are many fitting these few criteria. Perhaps one of them on your list is close?

OpenStudy (anonymous):

a. y=2x-8/x-3 b. y=2x^2 -8/x^2-9 c. y=2x^2-18/x^2-4 d. y=x^2-9/x^2-4

OpenStudy (tkhunny):

a) x = -3 is IN the Domain. No good. b) Pretty good. This is very close to what I have, excepting the "-8". This is it!! c) Both x = 2 and x = -2 are NOT in the Domain, but x = 3 and x = -3 are. No good. d) Horizontal asymptote is the x-axis. No good.